OpenStudy (anonymous):

Prove that this series diverges Summation {n=0)^infinity (2n)!/(n!)^2 Do you use the root or ratio test?

OpenStudy (turingtest):

ratio test works nicely

OpenStudy (anonymous):

i am pretty sure it does not does it help to know that \[\frac{(2n)!}{(n!)^2}=\binom{2n}{n}\]?

OpenStudy (anonymous):

oooh "diverges" of course the terms do not even go to zero

OpenStudy (anonymous):

does it matter if the index is n=0 and not n=1?

OpenStudy (anonymous):

makes no difference each term is \(\binom{2n}{n}\) so the terms get larger, not smaller

OpenStudy (anonymous):

I used the ratio test and got 0.. which im guessing is incorrect since i lost a point on my homework..

OpenStudy (anonymous):

yeah it is incorrect

OpenStudy (anonymous):

but the lim as n goest o inf. of 2n/n is indeterminable.. so you use lhopitals rule which gives you 2/1.. sooo?

OpenStudy (anonymous):

because 2>1 you knowit diverges?

OpenStudy (turingtest):

that's not quite what you get when you apply the ratio test, but even so the limit winds up being 2, yes

OpenStudy (anonymous):

when you apply the ratio test, you end up getting 1.. at least thats why i get when using the simplified (2n/n) that satellite just gave..

OpenStudy (turingtest):

\[{(2n+1)!\over(n+1)!} \cdot{n!\over (2n)!}=?\]

OpenStudy (anonymous):

shouldn't it be (2n+2)!/(n+1)!?

OpenStudy (anonymous):

aka (2(n+1))!/(n+1)!

OpenStudy (turingtest):

(n+1)!/n! = n+1

OpenStudy (anonymous):

2!/2! = 1 is what i get.. i factored out the 2! on both sides.. is that okay to do?

OpenStudy (turingtest):

no, you can't factor like that look at what I wrote above, you have to be able to cancel factors in factorial division (n+1)!/n! = n+1 think about why

OpenStudy (anonymous):

becasue (n+1)! = n!(n+1)

OpenStudy (turingtest):

exactly, so\[{(2n+1)!\over(n+1)!} \cdot{n!\over (2n)!}=\frac{n!}{(n+1)!}\cdot{(2n+1)!\over(2n)!}=? \]

OpenStudy (anonymous):

ok i think i got it.. the first term (n!/(n+1)!) = 1/(n+1) the second term (2(n+1))!/2n! is also equal to 1 because you make (2(n+1))! = (2n)!(n+1)/(2n)! which gives you (n+1) in the numerator so combining the two gives you 1?

OpenStudy (turingtest):

you can't pull the 2 out like that\[{(2n+1)!\over(n+1)!} \cdot{n!\over (2n)!}=\frac{n!}{(n+1)!}\cdot{(2n+1)!\over(2n)!}=\frac1{(n+1)}\cdot\frac{(2n+1)}1=\frac{2n+1}{n+1}\]

OpenStudy (turingtest):

if that confuses you, let 2n=x in the above

OpenStudy (turingtest):

\[{(x+1)!\over(n+1)!} \cdot{n!\over (x)!}=\frac{n!}{(n+1)!}\cdot{(x+1)!\over(x)!}=\frac1{(n+1)}\cdot\frac{(x+1)}1=\frac{x+1}{n+1}\]

OpenStudy (turingtest):

same rule, you just got confused because of the 2 :P

OpenStudy (anonymous):

if a_n is n!/(2n!) then a_(n+1) should be (n+1)!/(2(n+1))! this is probably where i'm getting most confused

OpenStudy (turingtest):

you are confusing (2n)!, the successor of which is (2n+1)!, with 2(n!), the successor of which is 2[(n+1)!]

OpenStudy (anonymous):

i thought if the parenthesis was around the entire term 2n then the factorial applies to both

OpenStudy (turingtest):

yes, it applies to both: imagine n=2, then (2n)!=4!=24 and when n=3 (2n)!=6!=720 but if it were the other way, when n=2 2(n!)=2(2!)=4 and when n=3 2(3!)=2(6)=12 see the difference?

OpenStudy (anonymous):

so when you are working out the root test.. and you have to add 1 to n.. you're really adding 1 to 2n since 2 is just a constant.. i think i understand better

OpenStudy (anonymous):

ratio test**

OpenStudy (anonymous):

thats why you told me to make 2n=x .. yay i understand it. ok great!

OpenStudy (turingtest):

Oh now I understand your point...actually, when you put it that way I realize I should be more careful; you can't just replace 2n with x for the reason you stated :/ however what satellite said is true, the factor of 2 will carry either way, though I owe it to you to do the root test properly. let me try that, sorry...

OpenStudy (turingtest):

yeah you can just distribute the 2, it's no big change...

OpenStudy (anonymous):

haha, ok ill wai until you're done to think more about this :D

OpenStudy (turingtest):

\[{(2(n+1))!\over(n+1)!} \cdot{n!\over (2n)!}={(2n+2)!\over(n+1)!} \cdot{n!\over (2n)!}=\\\frac1{(n+1)}\cdot\frac{(2n+2)(2n+1)}1=\frac{(2n+2)(2n+1)}{n+1}\]sorry, typos :P

OpenStudy (turingtest):

does that make sense?

OpenStudy (anonymous):

give me one second, im going tot ry to workit out the way you did

OpenStudy (anonymous):

where does the (2n+2)(2n+1) come from?

OpenStudy (anonymous):

oh nvm.. thats whats left after you divide (2n+2)!/2n! correct?

OpenStudy (turingtest):

right

OpenStudy (anonymous):

so the limit is infinity and it diverges.. that only took an hour haha thanks!!

OpenStudy (turingtest):

sorry I misunderstood your doubt earlier welcome!